KHW16 Item 13 The sound intensity 44.0 m from a wailing tornado siren is 0.150 W/m2 Part A What is the intensity at 1300 m ? Express your answer with the appropriate units alue Units Submit
Answer #1

Intensity is inversely proportional to the distance

I = k/r^2

r = 44 m

I = 0.150 w/m^2

k = I*r^2

k = 0.150*44^2 = 290.4 w

Therefore the intensity at r = 1300 m

I = 290.4/1300^2

I = 17 1.83 \small \mu w/m^2